OpenStudy (anonymous):

Determine the number of moles of compound in each of the following samples.? 4.50g H2O 471.6g Ba(OH)2 129.68g Fe3(PO4)2

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

do you know how to find the molar mass of a compound?

OpenStudy (unklerhaukus):

we can find the molar mass of a compound by adding up the molar masses of the atoms in it. you can find the molar mass of each type of atom using a periodic table

OpenStudy (unklerhaukus):

\[M_{\text H}=1.0 [\text {g/mole}]\]\[M_{\text O} =16.0[\text {g/mole}]\] \[M_{\text{H}_2\text{O}}=2\times M_{\text H}+M_{\text O}\]

OpenStudy (unklerhaukus):

once you have found the molar masses of the compounds divide the masses of the samples by the molar masses this will give you the number of moles in each sample

OpenStudy (unklerhaukus):

I'll give you an example: _____ Say you have 42.5 grams of Ammonia ( \(\text{NH}_3\) ) and you want to know the number of moles this is. •First we find the molar mass of Ammonia \[M_{\text{NH}_3}=M_{\text N}+3\times M_{\text H}\] using the periodic table we find \(M_{\text H}=1.0 [\text {g/mole}]\), \(M_{\text N}=14.0 [\text {g/mole}]\) \[M_{\text{NH}_3}=14.0[\text {g/mole}]+3\times 1.0[\text {g/mole}]\]\[\qquad\qquad=17.0[\text {g/mole}]\] •Now we divide mass by the molar mass to find the number of moles \[N_{\text{NH}_3}=\frac{m_{\text{NH}_3}}{M_{\text{NH}_3}}=\frac{42.5[\text g]}{17.0[\text {g/mole}]}=2.5[\text{mol}]\] ______ so we have found that 42.5 grams of ammonia is two and a half moles

OpenStudy (unklerhaukus):

@drasy22